OpenStudy (anonymous):

The conservation club washed 24 cars and earned $144 on Saturday. At this rate, how much money did they make on Sunday washing 30 cars? MEDAL,FAN,AND MESSAGE?

OpenStudy (anonymous):

proportion time... 24/144 = 30/x --- 24 cars to $144 = 30 cars to $x cross multiply (24)(x) = (30)(144) 24x = 4320 x = 4320/24 x = 180 They made $180 for washing 30 cars

OpenStudy (anonymous):

or you could do it this way.. 24 cars for 144.....144/24 = $6 per car 6 x 30 = 180 ($180 for 30 cars)

OpenStudy (anonymous):

thank you omg